OpenStudy (math2400):

Anyone get this? pic:

OpenStudy (math2400):

the 65-70 just if u could explain one that'd be helpful enough

OpenStudy (anonymous):

Just use your calculator and do \[65) \Theta = \sin^{-1}[-0.5640]\] note that for \[\sec(\Theta) = \frac{1}{\cos(\Theta)} = -1.116 \] So what you do is flip the terms. \[\cos(\Theta) = \frac{ 1 }{ -1.116 } = \Theta = \cos^{-1}[-0.8960]\]

OpenStudy (ash2326):

@Math2400 Are we allowed to use calculator?

OpenStudy (anonymous):

For the other ones similar to that you do the same thing .\[70) \csc(\Theta) = \frac{1}{\sin(\Theta)} = 1.485\]

OpenStudy (math2400):

yes i think so @ash2326. it's a friends and im half asleep and really dont feel like workin it out haha. Thanks guys!!

OpenStudy (math2400):

u guys dont need to do any others i got now

OpenStudy (anonymous):

Don't do your friends homework =P do mine plox.

OpenStudy (ash2326):

Just a min, if you use calculator you'll get only one solution. But each of these question wil have more than one solution

OpenStudy (math2400):

haha @mebs she just doesnt get it thought i'd help her. and really? so do u have to do it by hand then @ash2326

OpenStudy (ash2326):

We have to use calculator and hand both I'll show \[\sin x=-0.564\] Firstly use calculator to find \[x=\sin^{-1} (-0.564)\] This value will be in the range of [-90, 0] Now we have to give the answer in the range of (0, 360) DO you get this part?

OpenStudy (math2400):

ok i follow i believe

OpenStudy (ash2326):

Now tell me in which quadrants sin is negative?

OpenStudy (math2400):

III and IV

OpenStudy (ash2326):

Suppose using calculator we get -A so sin (-A)=-0.564 so sin A=0.564 Get this part?

OpenStudy (math2400):

ok yea u're just usin the -A as the number i got u now

OpenStudy (ash2326):

now the answers would be 180+A as sin(180+A)=-sin A and 360-A sin (360)=-sin A since sine is negative in these quadrats

OpenStudy (ash2326):

This is a bit tough part, Are you following?

OpenStudy (math2400):

yes im with u. u're doin great at explaining

OpenStudy (ash2326):

Thanks, so we're done the answers are 180+A and 360-A Now for the other questions, you have to keep in mind the quadrants in which they are positive or negative. It'd be better if you review inverse functions,

OpenStudy (math2400):

yea i got it. Thank u for explaining. Now let's see if i can explain it as well as u did to my friend haha.

OpenStudy (ash2326):

Cool :D